Please help me thank you so much!
A man who is 6 feet tall casts a shadow 8.4 feet long. A nearby building casts a shadow that is 42ft long. How high is the building?

Answer by seawind(8)

let x = the height of the building
6/8.4=x/42
8.4x=6*42
8.4x=252
x=252/8.4
x=30
the building is 30 feet.